OpenStudy (anonymous):

The graph of an equation is shown below: line joining ordered pairs negative 3,2 and 1, negative 2 Based on the graph, which of the following represents a solution to the equation? (–2,–3) (3, 1) (1, 3) (3, 2)

OpenStudy (anonymous):

@jhonyy9 @Luigi0210 @phi @SolomonZelman @robtobey @xapproachesinfinity

OpenStudy (anonymous):

i chose A but i got it wrong

OpenStudy (phi):

can you get in the habit of using a - rather than the word negative. The words make it harder to read (for me, at any rate)

OpenStudy (anonymous):

sorry thats how it comes up.

OpenStudy (phi):

They mention a graph.

OpenStudy (phi):

can you post the graph?

OpenStudy (phi):

to check the choice (-2,-3) put you finger on the x-axis (horizontal line) where it says -2. move your down to the line. how far did you move your finger? Count "boxes"

OpenStudy (phi):

ut you finger on the x-axis (horizontal line) where it says -2. move your down to the line. how far did you move your finger? Count "boxes" can you do that ?

OpenStudy (anonymous):

2 times

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

can you explain why i got it wrong

OpenStudy (phi):

Look at this graph. The purple line is less than a whole box down.

OpenStudy (phi):

That means when x is -2, the line is at (about) -.75 (just guessing... but not -1, which is too far down)

OpenStudy (phi):

If you put your finger at x=-3 you would go down to -2 but (-3,-2) is not any of your choices. Let's check (1,3)

OpenStudy (anonymous):

not that

OpenStudy (phi):

did you put your finger at x=1? now count boxes up until you reach the line. How many boxes do you count?

OpenStudy (anonymous):

2

OpenStudy (phi):

Here is what I see

OpenStudy (phi):

Is that what you did?

OpenStudy (anonymous):

no i counted up two boxes

OpenStudy (anonymous):

the point is in between the 4 and 2

OpenStudy (anonymous):

1,3 was the answer @phi

OpenStudy (phi):

Yes, 1,3 because if you put your finger at x=1 it takes "3 steps" (or 3 horizontal lines or 3 boxes) to get up to the line. the "y value" is 3 when x is 1

OpenStudy (anonymous):

well i know for next time

OpenStudy (anonymous):

can you explain this last question

OpenStudy (anonymous):

Two systems of equations are shown below: System A System B 6x + y = 2 2x – 3y = –10 –x - y = –3 –x – y = –3 Which of the following statements is correct about the two systems of equations? The value of x for System B will be 4 less than the value of x for System A because the coefficient of x in the first equation of System B is 4 less than the coefficient of x in the first equation of System A. They will have the same solution because the first equations of both the systems have the same graph. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 4 times the second equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –4 to the first equation of System A and the second equations are identical.